# MotionStreamer: Streaming Motion Generation via Diffusion-based Autoregressive Model in Causal Latent Space

### v2

@@ 2 Related Work
Text-conditioned motion generation.  Text-conditioned motion generation aims to synthesize 3D human motions from natural language descriptions [17]. Previous works [17, 39, 2] leverage VAE [27] to learn cross-modal mappings between text and motion spaces. Some works [53, 9, 12, 61, 4, 48, 58, 10, 3, 28, 15, 57, 25] also apply diffusion models [24] to this task. Another line of works [60, 26, 62, 65, 64, 33, 34] first discretize motions into discrete tokens and employ autoregressive models (e.g., GPT) for sequential token prediction. Furthermore, some approaches [19, 42, 41, 16, 37] adopt a BERT-style [13] bidirectional Transformer architecture [7] to reconstruct masked motion segments under text guidance. However, most existing works only focus on offline generation, where the entire motion sequence is generated at once. More recently, CAMDM [8] and AMDM [49] apply diffusion models into an autoregerssive manner for real-time interactive character control. Ready-to-React [6] further explores this idea in two-character interaction. CLoSD [54] and DART [63] apply it for real-time text driven motion control. However, these methods are not strictly causal, as they rely on a fixed-length context window while our method could handle variable-length historical information and incrementally generate motions in a streaming manner.
Motion compression.  Following the success of image generation [46], previous works [9, 60] first encode the raw motion sequences into a latent space and then generate motions within it. The most popular method is to use Vector Quantized Variational AutoEncoders (VQ-VAE) [56] for motion tokenization. TM2T [18] first introduces vector quantization for motion discretization. T2M-GPT [60] employs VQ-VAE to compress motion sequences into a discrete latent space and then uses a GPT for motion generation. MoMask [19] leverages Residual VQ-VAE (RVQ-VAE) [29] to progressively reduce quantization errors. In contrast, MLD [9] utilizes standard VAEs [27] to convert a motion sequence into an embedding and then use a diffusion model to generate the latent. However, the existing methods require a whole motion sequence to be encoded and decoded, which is not suitable for streaming generation. In this paper, we propose a causal motion compression approach to achieve streaming motion generation with online response.
@@ 3.1 Problem Formulation
Task definition.  We first introduce the formulation of streaming motion generation. In contrast to previous text-to-motion generation [17] which is conditioned on a pre-defined fixed text prompt, we consider the case where a series of text prompts are given sequentially. Given a sequence of text prompts $\{\mathcal{P}_{i}\}_{i=1}^{M}$, the goal is to generate a sequence of motion frames $\{x_{j}\}_{j=1}^{N}$, where $\mathcal{P}_{i}$ is the $i$-th text prompt and $x_{j}$ is the $j$-th frame pose.
Motion Representation.  Previous works [17, 26, 62, 19, 53] mainly uses the 263-dimensional pose representation [17] for motion generation. However, this representation requires an additional post-processing step [5], which is time-consuming and introduces rotation error [14] to be converted to SMPL [31] body parameters (see appendix Sec. E). To overcome this issue, we slightly modify it and directly use SMPL-based 6D rotation [31] as joint rotations. Similar to prior works on character control [51, 50], each pose $x$ is represented by a 272-dimensional vector:
@@
where we project the root on the XZ-plane (ground plane), $(\dot{r}^{x},\dot{r}^{z}\in\mathbb{R})$ are root linear velocities on the XZ-plane, $\dot{r}^{a}\in\mathbb{R}^{6}$ denotes root angular velocity represented in 6D rotations, $j^{p}\in\mathbb{R}^{3K}$, $j^{v}\in\mathbb{R}^{3K}$, and $j^{r}\in\mathbb{R}^{6K}$ are local joint positions, local velocities, and local rotations relative to the root space, $K$ is the number of joints. This representation removes the post-processing step and we could directly use it for animating a SMPL [31] character model.
@@ 3.2 Causal Temporal AutoEncoder
Architecture.  Causal TAE is designed to achieve continuous motion compression while explicitly modeling temporal dependencies and enforcing causal constraints for sequential motion representation. Fig. 3 shows the proposed Causal TAE network. We employ 1D causal convolution [59] for constructing temporal encoder $\mathcal{E}$ and decoder $\mathcal{D}$ to convert raw motion sequences into a causal latent space. The causality is guaranteed by a temporal padding scheme. Specifically, for a convolution layer with kernel size $k_{t}$, stride $s_{t}$ and dilation rate $d_{t}$, we pad $(k_{t}-1)\times d_{t}+(1-s_{t})$ frames at the beginning of the sequence. In this way, each frame only depends on the frames before it and the future frames are not involved in the computation. Moreover, explicitly modeling temporal causal structures in the latent space enables the model to learn temporal and causal dependencies inherent in the causally-related motion data more effectively.
Given a motion sequence $X=\{x_{1},x_{2},\cdots,x_{N}\}$ with $x_{t}\in\mathbb{R}^{D}$, where $N$ is the number of frames and $D$ is the motion dimension, we could obtain a set of temporal Gaussian distribution parameters $\{\mu_{1:N/l},\sigma^{2}_{1:N/l}\}$ and preform reparameterization [27] to get continuous motion latent representation $Z=\{z_{1},z_{2},\cdots,z_{N/l}\}$ with $z_{i}\in\mathbb{R}^{d_{c}}$, $l$ represents the temporal downsampling rate of the Encoder $\mathcal{E}$. This architecture reconstructs motion frames while strictly preserving temporal causality across the sequence.
@@ 3.3 MotionStreamer
Training.  Each training sample can be represented as: $S_{i}=(T_{i},C_{i},Z_{i})$, where $T_{i}\in\mathbb{R}^{1\times d_{t}}$ is the text embedding obtained via a pre-trained language model (e.g. T5-XXL [45]), $C_{i}\in\mathbb{R}^{k\times d_{c}}$ and $Z_{i}\in\mathbb{R}^{n\times d_{c}}$ are the historical motion latents and current motion latents encoded by the learned Causal TAE, where $k$, $n$, $d_{t}$ and $d_{c}$ denote the lengths of previous motion latents, current motion latents, historical motion latents, the text embedding dimension and the latent dimension respectively. We concatenate them along temporal axis to form a sequence $S_{i}=[T_{i},C_{i},Z_{i}]$. We employ a diffusion-based autoregressive Transformer to predict motion latents. The latent sequence $S$ is first processed by the Transformer and a causal mask is applied to ensure the temporal causality [60]. After the Transformer processing, we obtain the intermediate latents $\{c_{i}^{1},c_{i}^{2},\cdots,c_{i}^{n}\}$, which serve as the condition for the diffusion head (a small MLP) to predict motion latents $\{\hat{z}_{i}^{1},\hat{z}_{i}^{2},\cdots,End_{i}\}$. $End_{i}$ is the reference end latent inserted at the end of a sequence as the continuous stopping condition, which we will elaborate later. Following [30, 24], the loss function is defined as:
@@
Two-Forward strategy. We observe that using teacher-forcing [1] directly during training often leads to error accumulation in the autoregressive generation process. To this end, we propose a Two-Forward strategy that progressively introduces the test-time distribution during training. Specifically, after the first forward pass, we replace a subset of ground-truth motion latents with their generated counterparts, creating a mixture of real and generated motion latents. This hybrid input is then used in the second forward pass, where gradients are backpropagated to refine the model. We employ a cosine scheduler to control the proportion of replaced motion latents.
@@ 4.1 Experimental Setup
Dataset. We evaluate the proposed MotionStreamer on HumanML3D [17] and BABEL [43] datasets, with the original train and test splits. The HumanML3D dataset integrates motion sequences with three distinct textual descriptions. The BABEL dataset provides frame-level textual descriptions with explicit inter-segment transition labels. Unlike recent methods [48, 4] that use different motion representations for both HumanML3D and BABEL datasets, we employ the 272-dimensional motion representation as mentioned in Sec. 3.1 for both datasets. All motion sequences are uniformly resampled at 60 FPS.
@@ 4.2 Quantitative Results
Comparison on Text-to-Motion Generation. We trained an evaluator based on TMR [40] to evaluate the quality of the generated motions. We use the processed 272-dimensional motion data from HumanML3D [17] train set for text-to-motion model training. All baseline methods are trained from scratch following their original implementations. The comparison results on HumanML3D [17] test set are shown in Tab. 1. Our method outperforms baseline methods in multiple metrics.
Methods FID $\downarrow$ R@1 $\uparrow$ R@2 $\uparrow$ R@3 $\uparrow$ MM-D $\downarrow$ Div $\rightarrow$
Real motion 0.002 0.711 0.851 0.903 15.805 27.670
MDM [53] 22.557 0.524 0.693 0.773 17.223 27.355
MLD [9] 17.226 0.548 0.732 0.805 16.338 26.551
T2M-GPT [60] 11.175 0.608 0.772 0.831 16.810 27.617
MotionGPT [26] 14.175 0.436 0.598 0.668 17.890 27.014
MoMask [19] 10.731 0.622 0.782 0.850 16.128 27.317
AttT2M [64] 15.438 0.590 0.767 0.837 15.734 26.680
Ours 10.724 0.631 0.784 0.851 16.639 27.657
Table 1: Comparison with baseline text-to-motion generation methods on HumanML3D [17] test set. MM-D and Div denote Multimodal Distance and Diversity respectively.
Comparison on Long-Term Motion Generation. We adopt a Mix Training Strategy for streaming long-term generation training. Specifically, we create training samples by pairing adjacent subsequences from long motion sequences in BABEL. Additionally, we incorporate text-motion pairs from the HumanML3D dataset for mix training. The comparison results on BABEL [43] dataset are demonstrated in Tab. 2. We modified T2M-GPT to support streaming generation (marked as T2M-GPT*) and also adapted our model to use VQ for discretization (marked as VQ-LLaMA). Experimental results show that neither the existing long-term generation baseline nor the discrete autoregressive model performs as well as our streaming generation approach in the continuous latent space.
@@
[…]
@@
Table 2: Comparison with long-term motion generation methods on BABEL [43] dataset. Symbols $\uparrow$, $\downarrow$ and $\rightarrow$ indicate the higher, lower and closer to Ground Truth are better. Bold and underline indicate the best and second best results.
@@
Figure 4: Comparison on the First-frame Latency of different methods. The horizontal axis represents the number of generated frames, while the vertical axis indicates the time required to produce the first output frame.
@@ 4.4 Ablation Study
FID $\downarrow$ MPJPE $\downarrow$ FID $\downarrow$ R@3 $\uparrow$ MM-D. $\downarrow$ Div. $\rightarrow$
Real motion - - 0.002 0.903 15.805 27.670
VQ-VAE 5.173 63.9 11.024 0.834 16.792 27.614
AE 0.001 1.7 43.818 0.473 22.041 27.085
VAE 2.092 26.2 19.914 0.755 17.948 27.520
Ours 0.737 24.89 10.724 0.851 16.639 27.657
@@ 4.5 Applications
MotionStreamer offers various applications, including multi-round generation, long-term generation, and dynamic motion composition. (1) Multi-round generation requires iteratively generating motion in response to sequential or interactive textual inputs. Our model can process incremental text inputs, respond online, and autonomously determine when to stop generation. (2) Long-term generation requires smoothly generating long sequences of motion in response to sequential textual inputs. (3) Dynamic motion composition refers to the capability of seamlessly integrating diverse motion sequences while preserving the consistency of previously generated content as shown in Fig. 6. Our Causal TAE enables this application, and during the generation of subsequent motion latents, it eliminates the need for full-sequence re-decoding, requiring only the decoding of the newly generated latents. Please see the supplementary videos in our project page for more application demos.
AR Design choices FID $\downarrow$ R@3 $\uparrow$ MM-Dist $\downarrow$ Diversity $\rightarrow$
w/o QK Norm 11.127 0.839 16.525 27.530
w/o Two-Forward 11.978 0.847 16.440 27.703
w/o Diffusion Head 59.195 0.361 22.884 26.825
CLIP 14.033 0.792 17.564 27.328
Ours 10.724 0.851 16.639 27.657
@@ 5 Conclusion
We present MotionStreamer, a novel framework for streaming motion generation that integrates diffusion-based autoregressive model to directly predict causal motion latents. By introducing Causal TAE, MotionStreamer supports progressive textual inputs and enables online response. To mitigate cumulative errors in the streaming generation process, we propose a Two-Forward training strategy. Our method outperforms baseline approaches, demonstrating its competitiveness in motion generation while providing greater flexibility. It can be applied to multi-round generation, long-term generation, and dynamic motion composition. Future work could explore hybrid strategies that allow bidirectional refinement without compromising streaming generation. One potential way is to predict a set of future latents at each step, which could enable motion in-between and localized editing while preserving streaming manner.
Limitations.  Despite its effectiveness, the streaming generation paradigm limits the applications of motion in-betweening and localized editing of intermediate tokens, as it inherently relies on unidirectional modeling. This limitation restricts flexibility in scenarios requiring fine-grained adjustments, such as seamlessly inserting new motions between existing frames or interactively refining motion details while preserving global coherence.
Appendix

### v3

@@ 2 Related Work
Text-conditioned motion generation.  Text-conditioned motion generation aims to synthesize 3D human motions from natural language descriptions [20]. Various approaches [45, 71, 35, 8, 73, 75, 29, 51, 44, 50, 9, 49, 69, 70, 79, 10, 28] have explored diverse architectures and methodologies to enhance the naturalness and expressiveness for digital human and motion animation. Previous works [20, 47, 2] leverage VAE [32] to learn cross-modal mappings between text and motion spaces. Some works [64, 12, 14, 77, 4, 59, 72, 3, 33, 17, 68, 30] also apply diffusion models [27] to this task. Another line of works [76, 31, 78, 82, 81, 39, 40, 19] first discretize motions into discrete tokens and employ autoregressive models (e.g., GPT) for sequential token prediction. Furthermore, some approaches [22, 53, 52, 18, 43] adopt a BERT-style [15] bidirectional Transformer architecture [7] to reconstruct masked motion segments under text guidance. However, most existing works only focus on offline generation, where the entire motion sequence is generated at once. More recently, CAMDM [11] and AMDM [60] apply diffusion models into an autoregerssive manner for real-time interactive character control. Ready-to-React [6] further explores this idea in two-character interaction. CLoSD [65] and DART [80] apply it for real-time text driven motion control. However, these methods are not strictly causal, as they rely on a fixed-length context window while our method could handle variable-length historical information and incrementally generate motions in a streaming manner.
Motion compression.  Following the success of image generation [57], previous works [12, 76] first encode the raw motion sequences into a latent space and then generate motions within it. The most popular method is to use Vector Quantized Variational AutoEncoders (VQ-VAE) [67] for motion tokenization. TM2T [21] first introduces vector quantization for motion discretization. T2M-GPT [76] employs VQ-VAE to compress motion sequences into a discrete latent space and then uses a GPT for motion generation. MoMask [22] leverages Residual VQ-VAE (RVQ-VAE) [34] to progressively reduce quantization errors. In contrast, MLD [12] utilizes standard VAEs [32] to convert a motion sequence into an embedding and then use a diffusion model to generate the latent. However, existing methods require a whole motion sequence to be encoded and decoded, which is not suitable for streaming generation. In this paper, we propose a causal motion compression approach to achieve streaming motion generation with online response.
@@ 3.1 Problem Formulation
Task definition.  We first introduce the formulation of streaming motion generation. In contrast to previous text-to-motion generation [20] which is conditioned on a pre-defined fixed text prompt, we consider the case where a series of text prompts are given sequentially. Given a streaming sequence of text prompts $\{\mathcal{P}_{i}\}_{i=1}^{M}$, the goal is to generate a sequence of motion frames $\{x_{j}\}_{j=1}^{N}$ online, where $\mathcal{P}_{i}$ is the $i$-th text prompt and $x_{j}$ is the $j$-th frame pose.
Motion Representation.  Previous works [20, 31, 78, 22, 64] mainly uses the 263-dimensional pose representation [20] for motion generation. However, this representation requires an additional post-processing step [5], which is time-consuming and introduces rotation error [16] to be converted to SMPL [37] body parameters. To overcome this issue, we slightly modify it and directly use SMPL-based 6D rotation [37] as joint rotations. Similar to prior works on character control [62, 61], each pose $x$ is represented by a 272-dimensional vector:
@@
where we project the root on the XZ-plane (ground plane), $(\dot{r}^{x},\dot{r}^{z}\in\mathbb{R})$ are root linear velocities on the XZ-plane, $\dot{r}^{a}\in\mathbb{R}^{6}$ denotes root angular velocity represented in 6D rotations, $j^{p}\in\mathbb{R}^{3K}$, $j^{v}\in\mathbb{R}^{3K}$, and $j^{r}\in\mathbb{R}^{6K}$ are local joint positions, local velocities, and local rotations relative to the root space, $K$ is the number of joints. For SMPL [37] character, $K=22$, then we get the $2+6+3\times 22+3\times 22+6\times 22=272$ dimensions. This representation removes the post-processing step and we could directly use it for animating a SMPL character model.
@@ 3.2 Causal Temporal AutoEncoder
Architecture.  Causal TAE is designed to achieve continuous motion compression while explicitly modeling temporal dependencies and enforcing causal constraints for sequential motion representation. Fig. 3 shows the proposed Causal TAE network. We employ 1D causal convolution [74] for constructing temporal encoder $\mathcal{E}$ and decoder $\mathcal{D}$ to convert raw motion sequences into a causal latent space. The causality is guaranteed by a temporal padding scheme. Specifically, for a convolution layer with kernel size $k_{t}$, stride $s_{t}$ and dilation rate $d_{t}$, we pad $(k_{t}-1)\times d_{t}+(1-s_{t})$ frames at the beginning of the sequence. In this way, each frame only depends on the frames before it and the future frames are not involved in the computation. Moreover, explicitly modeling temporal causal structures in the latent space enables the model to learn temporal and causal dependencies inherent in the causally-related motion data.
Given a motion sequence $X=\{x_{1},x_{2},\cdots,x_{N}\}$ with $x_{t}\in\mathbb{R}^{D}$, where $N$ is the number of frames and $D$ is the motion dimension ($D=272$), we could obtain a set of temporal Gaussian distribution parameters $\{\mu_{1:N/l},\sigma^{2}_{1:N/l}\}$ and preform reparameterization [32] to get continuous motion latent representation $Z=\{z_{1},z_{2},\cdots,z_{N/l}\}$ with $z_{i}\in\mathbb{R}^{d_{c}}$, $l$ represents the temporal downsampling rate of the Encoder $\mathcal{E}$. This architecture reconstructs motion frames while strictly preserving temporal causality across the sequence.
@@ 3.3 MotionStreamer
Training.  Each training sample can be represented as: $S_{i}=(T_{i},C_{i},Z_{i})$, where $T_{i}\in\mathbb{R}^{1\times d_{t}}$ is the text embedding obtained via a pre-trained language model (e.g. T5-XXL [56]), $C_{i}\in\mathbb{R}^{k\times d_{c}}$ and $Z_{i}\in\mathbb{R}^{n\times d_{c}}$ are the previous motion latents and current motion latents encoded by the learned Causal TAE, where $k$, $n$, $d_{t}$ and $d_{c}$ denote the lengths of previous motion latents, current motion latents, the text embedding dimension and the latent dimension respectively. We concatenate them along temporal axis to form a sequence $S_{i}=[T_{i},C_{i},Z_{i}]$. We employ a diffusion-based autoregressive Transformer to predict motion latents. The latent sequence $S$ is first processed by the Transformer and a causal mask is applied to ensure the temporal causality [76]. After the Transformer processing, we obtain the intermediate latents $\{c_{i}^{1},c_{i}^{2},\cdots,c_{i}^{n}\}$, which serve as the condition for the diffusion head (a small MLP) to predict motion latents $\{\hat{z}_{i}^{1},\hat{z}_{i}^{2},\cdots,End_{i}\}$. $End_{i}$ is the reference end latent inserted at the end of a sequence as the continuous stopping condition, which we will elaborate later. Following [36, 27], the loss function is defined as:
@@
Two-Forward strategy. We observe that using teacher-forcing [1] directly during training often leads to error accumulation in the autoregressive generation process. To this end, we propose a Two-Forward strategy that progressively introduces the test-time distribution during training. Specifically, after the first forward pass, we replace a subset of ground-truth motion latents with their generated counterparts, creating a mixture of real and generated motion latents. This hybrid input is then used in the second forward pass, where gradients are backpropagated. We employ a cosine scheduler to control the proportion of replaced motion latents. More details are provided in Sec.A of the appendix.
@@ 4.1 Experimental Setup
Dataset. We evaluate the proposed MotionStreamer on HumanML3D [20] and BABEL [54] datasets, with the original train and test splits. The HumanML3D dataset integrates motion sequences with three distinct textual descriptions. The BABEL dataset provides frame-level textual descriptions with explicit inter-segment transition labels. Unlike recent methods [59, 4] that use different motion representations for both HumanML3D and BABEL datasets, we employ the 272-dimensional motion representation as mentioned in Sec. 3.1 for both datasets. All motion sequences are uniformly resampled to 30 FPS.
@@ 4.2 Quantitative Results
Comparison on Text-to-Motion Generation. We trained an evaluator based on TMR [48] to evaluate the quality of the generated motions. We use the processed 272-dimensional motion data from HumanML3D [20] train set for text-to-motion model training. All baselines are trained from scratch following their original implementations. As shown in Tab. 1, our method receives better performance across multiple metrics on HumanML3D [20] test set .
Methods FID $\downarrow$ R@1 $\uparrow$ R@2 $\uparrow$ R@3 $\uparrow$ $M$-D $\downarrow$ Div $\rightarrow$
Real motion 0.002 0.702 0.864 0.914 15.151 27.492
MDM [64] 23.454 0.523 0.692 0.764 17.423 26.325
MLD [12] 18.236 0.546 0.730 0.792 16.638 26.352
T2M-GPT [76] 12.475 0.606 0.774 0.838 16.812 27.275
MotionGPT [31] 14.375 0.456 0.598 0.628 17.892 27.114
MoMask [22] 12.232 0.621 0.784 0.846 16.138 27.127
AttT2M [81] 15.428 0.592 0.765 0.834 15.726 26.674
Ours 11.790 0.631 0.802 0.859 16.081 27.284
Table 1: Comparison with baseline text-to-motion generation methods on HumanML3D [20] test set. $M$-D and Div denote Multimodal Distance and Diversity respectively.
Comparison on Long-Term Motion Generation. We adopt a Mix Training Strategy for streaming long-term generation training. Specifically, we create training samples by pairing adjacent subsequences from long motion sequences in BABEL. Additionally, we incorporate text-motion pairs from the HumanML3D dataset for mix training. The comparison results on BABEL [54] dataset are demonstrated in Tab. 2. Following FlowMDM [4], the motion transition length is set to 30 frames. We modified T2M-GPT to support streaming generation (marked as T2M-GPT*) and also adapted our model to use VQ for discretization (marked as VQ-LLaMA). Experimental results show that neither the existing long-term generation baseline nor the discrete autoregressive model performs as well as our streaming generation approach in the continuous latent space.
@@
R@3 $\uparrow{}$ FID $\downarrow{}$ Div $\rightarrow$ $M$-D $\downarrow{}$ FID $\downarrow{}$ Div $\rightarrow{}$ PJ $\rightarrow{}$ AUJ $\downarrow{}$
@@
Table 2: Comparison with long-term motion generation methods on BABEL [54] dataset.
@@
Figure 4: Comparison on the First-frame Latency of different methods. Horizontal axis: the number of generated frames. Vertical axis: the time required to produce the first output frame.
@@ 4.4 Ablation Study
FID $\downarrow$ MPJPE $\downarrow$ FID $\downarrow$ R@3 $\uparrow$ $M$-D $\downarrow$ Div $\rightarrow$
Real motion - - 0.002 0.914 15.151 27.492
VQ-VAE 5.173 63.9 13.226 0.824 16.746 27.024
AE 0.001 1.7 43.828 0.463 22.040 27.382
VAE 2.092 26.2 19.902 0.735 17.926 27.312
Ours 0.661 22.9 11.790 0.859 16.081 27.284
@@ 4.5 Applications
MotionStreamer offers various applications, including multi-round generation, long-term generation, and dynamic motion composition. (1) Multi-round generation requires iteratively generating motion in response to sequential or interactive textual inputs. Our model can process incremental text inputs, respond online, and autonomously determine when to stop generation. (2) Long-term generation requires smoothly generating long motion sequences in response to sequential textual inputs. (3) Dynamic motion composition refers to the capability of seamlessly integrating diverse motion sequences while preserving the consistency of previously generated motion prefix, as shown in Fig. 6. Our Causal TAE enables this application, which eliminates the need for full-sequence re-decoding during the generation of subsequent motion latents, requiring only the decoding of the newly generated motion latents.
AR Design choices FID $\downarrow$ R@3 $\uparrow$ $M$-D $\downarrow$ Div $\rightarrow$
Real motion 0.002 0.914 15.151 27.492
w/o QK Norm 12.324 0.836 16.524 27.126
w/o Two-Forward 12.944 0.842 16.432 27.282
w/o Diffusion Head 59.296 0.360 22.864 27.325
CLIP 14.234 0.791 17.534 27.316
Ours 11.790 0.859 16.081 27.284
@@ 5 Conclusion
We present MotionStreamer, a novel framework for streaming motion generation that integrates diffusion-based autoregressive model to directly predict causal motion latents. By introducing Causal TAE, MotionStreamer supports online response with progressive textual inputs. We propose a Two-Forward training strategy to mitigate cumulative errors in streaming generation process. Our method outperforms baseline approaches, demonstrating its competitiveness in motion generation while providing greater flexibility. It can be applied to multi-round motion generation, long-term motion generation, and dynamic motion composition.
Acknowledgement
This work was partially supported by the National Key R&D Program of China (No. 2024YFB2809102), NSFC (No. 62402427, NO. U24B20154), Zhejiang Provincial Natural Science Foundation of China (No. LR25F020003), DeepGlint, Zhejiang University Education Foundation Qizhen Scholar Foundation, and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University, the National Key R&D Program of China (2022ZD0160201), and Shanghai Artificial Intelligence Laboratory.
Appendix of “MotionStreamer: Streaming Motion Generation via Diffusion-based Autoregressive Model in Causal Latent Space”